# Context-free graphs and their transition groups

### v2

@@ 8. Rationality of CF-TR
An asynchronous transducer is a quintuple $T=(Q,q_{0},\Sigma,\delta,o)$, where
@@
For the sake of brevity, we denote $\delta(q,\sigma)=\widetilde{q}$ and $o(q,\sigma)=w$ simply with $q\smash{\stackrel{{\scriptstyle\text{\tiny{$\sigma|w$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{q}$. It is straightforward that one can extend $o$ and $\delta$ to finite and infinite words. More precisely, if $\widehat{\Sigma}=\Sigma^{*}\cup\Sigma^{\omega}$, we have $\widehat{o}:Q\times\widehat{\Sigma}\rightarrow\widehat{\Sigma}$ and $\widehat{\delta}:Q\times\widehat{\Sigma}\rightarrow Q$ simply by the formula
@@
A homeomorphism $\varphi:\Sigma^{\omega}\rightarrow\Sigma^{\omega}$ is rational if there exists an asynchronous transducer $(Q,q_{0},\Sigma,\delta,o)$ such that $\varphi(\eta)=q_{0}\ast\eta$ for all $\eta\in\Sigma^{\omega}$. In [14], it has been proved that the set of rational homeomorphisms over an alphabet $\Sigma$ is indeed a group $\mathcal{R}_{\Sigma}$ (called the rational group) and any two rational groups over two different alphabets are isomorphic (so it makes sense to simply write $\mathcal{R}$). In [4] it is shown that the rational group $\mathcal{R}$ is simple and non-finitely generated. We say that a group which embeds in any rational group is rational. From a topological point of view $\Sigma^{\omega}$ is a Cantor space where the metric is given by the longest common prefix. In this way, there is a well-defined notion of bi-lipschitz homeomorphism. A transducer associated to a homeomorphism is said to be bi-lipschitz if the homeomorphism is bi-lipschtz. The family of groups which can be represented by bi-lipschitz transducers includes Thompson-like groups, self-similar groups and hyperbolic groups (see [3]). We want to construct a bi-lipschitz transducer for each element of our group belonging to CF-TR, so that any such group embeds in $\mathcal{R}$. To fix the notation, let $\{\Gamma_{0},\ldots,\Gamma_{N}\}$ be the representative of each end-isomorphim class of the context-free inverse graph $(\Gamma,x_{0})$, and let $\{\Delta_{0},\ldots,\Delta_{N}\}$ be the corresponding frontier vertices. We put $\Gamma_{0}=\Gamma$. As we have discussed at the end of Lemma 5.1, for each end-cone $\Gamma(x,x_{0})$ of a context-free graph inverse graph $(\Gamma,x_{0})$ we have a certain number of second-level end-cones which are end-isomorphic to some representative among $\{\Gamma_{0},\ldots,\Gamma_{N}\}$. This means that each $\Gamma_{i}$ has $k(i)$ second-level end-cones that we denote by $\Gamma^{(i)}_{1},\ldots,\Gamma^{(i)}_{k(i)}$. Each subgraph $\Gamma^{(i)}_{j}$ is end isomorphic to some end-cone, say $\Gamma_{h}\in\{\Gamma_{0},\ldots,\Gamma_{n}\}$, and let us denote by $\Psi(i,j):\Gamma^{(i)}_{j}\to\Gamma_{h}$ such end-isomorphism and put $\psi(i,j)=h$, where $i\in[0,N]$, $j\in[1,k(i)]$. We also denote by $\Delta_{j}^{(i)}$ the frontier vertices of $\Gamma^{(i)}_{j}$. We now define the alphabet on which our transducer will act:
@@
The map $\Xi$
$\Xi:V(\Gamma)\to\{w\Lambda^{\omega}:w\in\mathcal{F}(\Lambda)\}$
from the vertices of $\Gamma$ to the set of infinite words starting with a well-formed word defined by $\Xi(x)=w[x]\Lambda^{\omega}$ is a bijection.
This is a key point that will be used in proving that the action of each letter $a\in A$ on $\Gamma$ is rational. We now define the asynchronous transducer $\mathcal{T}_{\Gamma}$. This transducer acts on $\Lambda$ and has set of states
@@
We now describe the set of transitions. The state $E$ is a sink state, which means that we have the transitions $E\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda\mid\lambda$}}}}{{\relbar\joinrel\longrightarrow}}}E$ for all $\lambda\in\Lambda$. For each fixed $a\in A$ we now describe a transducer that mimics the action of the letter $a$ when we are in a vertex $x$ encoded in the closed set $\Xi(x)$. This is done by first checking whether the word in the input encodes a well-formed word representing a vertex $x$, and then changing the last letters to match the new vertex $x\cdot a$ in the new encoding in the output. The transitions are defined as follows. We start by describing how the initial state acts:
(4) $\begin{cases}(\Gamma_{0},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|1$}}}}{{\relbar\joinrel\longrightarrow}}}(\Gamma_{0},a)&\mbox{if }\lambda=(\Gamma_{0})\\ (\Gamma_{0},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda$}}}}{{\relbar\joinrel\longrightarrow}}}E&\mbox{ in the other cases when }\lambda\mbox{ is not final. }\end{cases}$
@@
(5) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}(\lambda(1),a),\mbox{ where }\lambda^{\prime}=(\Gamma_{j}^{(i)}),\mbox{\bf if }\lambda(2)=0\mbox{ and }\lambda(1)<\Gamma_{j}^{(i)};\$
@@
$\Gamma_{t}^{(m)}<\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)},\mbox{ and }\Psi(s,h)(\lambda(2))\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z\mbox{ is a walk in }\Psi(s,h)(\Gamma_{s}^{(h)})=\Gamma_{m},\mbox{ with }z\in\Delta_{t}^{(m)}.$
@@
We recall that for a vertex $x\in\Gamma$ and a letter $a\in A$ we denote by $x\cdot a$ the vertex such that $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}(x\cdot a)$ is an edge in $\Gamma$. We have the following lemma.
@@
  • Suppose that condition (6) holds. In this case $v_{\ell}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\Psi(i,j)(z)$ is an edge of $\Gamma_{h}$ with $z\in\Delta_{j}^{(i)}$, and so by (10) and by the definition we get:
@@
    We claim that $\Xi(x\cdot a)=\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)\Lambda^{\omega}$. Now, since $\Psi(i,j)^{-1}(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j}^{(i)}$ and $\varphi_{\ell}(x)=\varphi_{\ell-1}(x)$, we deduce that also $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell-1}(z)$ is an edge in $\Gamma(x_{\ell-1},x_{0})$ with $x\cdot a=\varphi^{-1}_{\ell-1}(z)\in\Delta(x_{\ell-1},x_{0})$. Therefore, by Lemma 8.1, the well-formed word $\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)$ is associated the rooted end-geodesic:
@@
    where $v_{\ell}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j_{\ell}}^{(i_{\ell})}$ with $z\in\Delta_{j_{\ell}}^{(i_{\ell})}$. Now, since $\varphi_{\ell}(v_{\ell})=x$ and $\varphi^{-1}_{\ell}(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell}(z)$ is an edge in $\Gamma(x_{\ell},x_{0})$, we conclude that $x\cdot a=\varphi^{-1}_{\ell}(z)$. Hence, $\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})},z)$ is the well-formed word associated to the rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell-1},x_{0})\supset\Gamma(x\cdot a,x_{0})$, i.e., $(\Gamma_{0},a)\ast w[x]\eta=w[x\cdot a]\eta$ for all $\eta\in\Lambda^{\omega}$.
@@
    where $\Gamma_{t}^{(m)}<\Gamma_{s}^{(h)}=\Gamma_{j_{\ell}}^{(i_{\ell})}$ and $\Psi(s,h)(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge in $\Gamma_{m}$ with $z\in\Delta_{t}^{(m)}$. Since $\Gamma_{t}^{(m)}$ is a second-level end-cone of $\Gamma_{m}$ and $\phi=\Psi(s,h)\circ\varphi_{\ell}:\Gamma(x,x_{0})\to\Gamma_{m}$ is an end-isomorphism, we deduce that $\phi^{-1}(\Gamma_{t}^{(m)})$ is a second level end-cone of $\Gamma(x,x_{0})$ containing $\phi^{-1}(z)$, and since $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\phi^{-1}(z)$ is an edge in $\Gamma(x,x_{0})$ we have $\phi^{-1}(\Gamma_{t}^{(m)})=\Gamma(x\cdot a,x_{0})$. Take $\varphi_{\ell+1}=\left.\phi\right|_{\Gamma(x\cdot a,x_{0})}$. We clearly have $(\left.\Psi(s,h)\circ\varphi_{\ell})\right|_{\Gamma(x\cdot a,x_{0})}=\varphi_{\ell+1}$. Thus, by Lemma 8.1 we have that the rooted end-geodesic
@@
The transition group $\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$ of a collection of context-free complete graphs $\Gamma^{(1)}\ldots,\Gamma^{(k)}$ is rational.
@@
As a shorthand notation let us put $\gamma_{a}=(\Gamma_{0},a)$, $a\in A$. By Lemma 8.5 and the fact that $\Gamma$ is an inverse graph, we have $(\gamma_{a^{-1}}\gamma_{a})\ast w[x]\eta=w[x]\eta$ and $(\gamma_{a}\gamma_{a^{-1}})\ast w[x]\eta=w[x]\eta$ for all $\eta\in\Lambda^{\omega}$. Moreover, by Lemma 8.4 for all $\xi\in\Lambda^{\omega}\setminus\mathcal{F}(\Lambda)$ we have $(\gamma_{a^{-1}}\gamma_{a})\ast\xi=\xi$ and $(\gamma_{a}\gamma_{a^{-1}})\ast\xi=\xi$. Therefore, we may conclude that each state $\gamma_{a}$ of $\mathcal{T}_{\Gamma}$ gives rise to a rational bijection $\gamma_{a}\ast:\Lambda^{\omega}\to\Lambda^{\omega}$ with inverse $\gamma_{a^{-1}}\ast$. Thus, by [3, Proposition 1.3], each state $\gamma_{a}$ is a rational homeomorphism of $\Lambda^{\omega}$. Let us prove that the subgroup $H(\Gamma)=\langle\gamma_{a}:a\in A\rangle$ of the rational group $\mathcal{R}_{\Lambda}$ of all the rational homeomorphisms of $\Lambda^{\omega}$ is isomorphic to the transtion group $\mathcal{G}(\Gamma)$. Consider a word $u=a_{1}\ldots a_{p}\in A^{*}$. By Lemma 8.5, and using a simple induction on the length $p$ of a word $u$ we have that the following equality
@@
9. Free product of context-free inverse graphs

### v3

@@ 7. Rationality of CF-TR
Before turning to the questions discussed in the previous section, we first examine a property common to all currently known examples of CoCF groups and shared by our class. A sequential transducer is a quintuple $T=(Q,q_{0},\Sigma,\delta,o)$, where
@@
See, for instance, [3]. For the sake of brevity, we denote $\delta(q,\sigma)=\widetilde{q}$ and $o(q,\sigma)=w$ simply with the diagram $q\smash{\stackrel{{\scriptstyle\text{\tiny{$\sigma|w$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{q}$. It is straightforward that one can extend $o$ and $\delta$ to finite and infinite words. More precisely, if $\widehat{\Sigma}=\Sigma^{*}\cup\Sigma^{\omega}$, we have $\widehat{o}:Q\times\widehat{\Sigma}\rightarrow\widehat{\Sigma}$ and $\widehat{\delta}:Q\times\widehat{\Sigma}\rightarrow Q$ simply by the formula
@@
A homeomorphism $\varphi:\Sigma^{\omega}\rightarrow\Sigma^{\omega}$ is rational if there exists a sequential transducer $(Q,q_{0},\Sigma,\delta,o)$ such that $\varphi(\eta)=q_{0}\ast\eta$ for all $\eta\in\Sigma^{\omega}$.
Definition 7.1 ([15]).
The set of rational homeomorphisms over an alphabet $\Sigma$ is a group $\mathcal{R}_{\Sigma}$ (called the rational group).
Any two rational groups over two different alphabets are isomorphic (so it makes sense to simply write $\mathcal{R}$).
In [5] it is shown that the rational group $\mathcal{R}$ is simple and non-finitely generated. We say that a group which embeds in any rational group is rational. From a topological point of view $\Sigma^{\omega}$ is a Cantor space where the metric is given by the longest common prefix. In this way, there is a well-defined notion of bi-lipschitz homeomorphism. A transducer associated to a homeomorphism is said to be bi-lipschitz if the homeomorphism is bi-lipschtz. The family of groups which can be represented by bi-lipschitz transducers includes Thompson-like groups, self-similar groups and hyperbolic groups (see [4]). We want to construct a bi-lipschitz transducer for each element of our group belonging to CF-TR, so that any such group embeds in $\mathcal{R}$. To fix the notation, let $\{\Gamma_{0},\ldots,\Gamma_{N}\}$ be the representative of each end-isomorphim class of the context-free graph $(\Gamma,x_{0})$, and let $\{\Delta_{0},\ldots,\Delta_{N}\}$ be the corresponding frontier vertices. We put $\Gamma_{0}=\Gamma$. As we have discussed at the end of Lemma 5.1, for each end-cone $\Gamma(x,x_{0})$ of a context-free graph graph $(\Gamma,x_{0})$ we have a certain number of second-level end-cones which are end-isomorphic to some representative among $\{\Gamma_{0},\ldots,\Gamma_{N}\}$. This means that each $\Gamma_{i}$ has $k(i)$ second-level end-cones that we denote by $\Gamma^{(i)}_{1},\ldots,\Gamma^{(i)}_{k(i)}$. Each subgraph $\Gamma^{(i)}_{j}$ is end isomorphic to some end-cone, say $\Gamma_{h}\in\{\Gamma_{0},\ldots,\Gamma_{n}\}$, and let us denote by $\Psi(i,j):\Gamma^{(i)}_{j}\to\Gamma_{h}$ such end-isomorphism and put $\psi(i,j)=h$, where $i\in[0,N]$, $j\in[1,k(i)]$. We also denote by $\Delta_{j}^{(i)}$ the frontier vertices of $\Gamma^{(i)}_{j}$. We now define the alphabet on which our transducer will act:
@@
The map $\Xi:V(\Gamma)\to\mathcal{F}(\Lambda)\Lambda^{\omega}$, defined by $\Xi(x)=w[x]\Lambda^{\omega}$, $\forall x\in V(\Gamma)$, is a bijection.
This is a key point that will be used in proving that the action of each letter $a\in A$ on $\Gamma$ is rational. We now define the sequential transducer $\mathcal{T}_{\Gamma}$. This transducer acts on $\Lambda$ and has set of states
@@
We now describe the set of transitions. The state $E$ is a sink state, which means that we have the transitions $E\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$\lambda\mid\lambda$}}}{{\relbar\joinrel\longrightarrow}}}E$ for all $\lambda\in\Lambda$. For each fixed $a\in A$ we now describe a transducer that mimics the action of the letter $a$ when we are in a vertex $x$ encoded in the closed set $\Xi(x)$. This is done by first checking whether the word in the input encodes a well-formed word representing a vertex $x$, and then changing the last letters to match the new vertex $x\cdot a$ in the new encoding in the output. The transitions are defined as follows. We start by describing how the initial state acts:
[…]
@@
(4) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}(\lambda(1),a),\mbox{ where }\lambda^{\prime}=(\Gamma_{j}^{(i)}),\mbox{\bf if }\lambda(2)=0\mbox{ and }\lambda(1)<\Gamma_{j}^{(i)};\\$
@@
[…]
@@
We recall that for a vertex $x\in\Gamma$ and a letter $a\in A$ we denote by $x\cdot a$ the vertex such that $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}(x\cdot a)$ is an edge in $\Gamma$. We have the following lemma.
@@
  • Suppose that condition (5) holds. In this case $v_{\ell}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\Psi(i,j)(z)$ is an edge of $\Gamma_{h}$ with $z\in\Delta_{j}^{(i)}$, and so by (9) and by the definition we get:
@@
    We claim that $\Xi(x\cdot a)=\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)\Lambda^{\omega}$. Now, since $\Psi(i,j)^{-1}(v_{\ell})\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j}^{(i)}$ and $\varphi_{\ell}(x)=\varphi_{\ell-1}(x)$, we deduce that also $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell-1}(z)$ is an edge in $\Gamma(x_{\ell-1},x_{0})$ with $x\cdot a=\varphi^{-1}_{\ell-1}(z)\in\Delta(x_{\ell-1},x_{0})$. Therefore, by Lemma 7.2, the well-formed word $\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)$ is associated the rooted end-geodesic:
@@
    where $v_{\ell}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j_{\ell}}^{(i_{\ell})}$ with $z\in\Delta_{j_{\ell}}^{(i_{\ell})}$. Now, since $\varphi_{\ell}(v_{\ell})=x$ and $\varphi^{-1}_{\ell}(v_{\ell})\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell}(z)$ is an edge in $\Gamma(x_{\ell},x_{0})$, we conclude that $x\cdot a=\varphi^{-1}_{\ell}(z)$. Hence, $\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})},z)$ is the well-formed word associated to the rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell-1},x_{0})\supset\Gamma(x\cdot a,x_{0})$, i.e., $(\Gamma_{0},a)\ast w[x]\eta=w[x\cdot a]\eta$ for all $\eta\in\Lambda^{\omega}$.
@@
    where $\Gamma_{t}^{(m)}<\Gamma_{s}^{(h)}=\Gamma_{j_{\ell}}^{(i_{\ell})}$ and $\Psi(s,h)(v_{\ell})\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge in $\Gamma_{m}$ with $z\in\Delta_{t}^{(m)}$. Since $\Gamma_{t}^{(m)}$ is a second-level end-cone of $\Gamma_{m}$ and $\phi=\Psi(s,h)\circ\varphi_{\ell}:\Gamma(x,x_{0})\to\Gamma_{m}$ is an end-isomorphism, we deduce that $\phi^{-1}(\Gamma_{t}^{(m)})$ is a second level end-cone of $\Gamma(x,x_{0})$ containing $\phi^{-1}(z)$, and since $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\phi^{-1}(z)$ is an edge in $\Gamma(x,x_{0})$ we have $\phi^{-1}(\Gamma_{t}^{(m)})=\Gamma(x\cdot a,x_{0})$. Take $\varphi_{\ell+1}=\left.\phi\right|_{\Gamma(x\cdot a,x_{0})}$. We clearly have $(\left.\Psi(s,h)\circ\varphi_{\ell})\right|_{\Gamma(x\cdot a,x_{0})}=\varphi_{\ell+1}$. Thus, by Lemma 7.2 we have that the rooted end-geodesic
@@
The transition group $\mathcal{G}\bigl(\Gamma^{(1)}\sqcup\cdots\sqcup\Gamma^{(k)}\bigr)$ of a collection of context-free complete graphs $\Gamma^{(1)}\ldots,\Gamma^{(k)}$ is rational.
@@
As a shorthand notation let us put $\gamma_{a}=(\Gamma_{0},a)$, $a\in A$. By Lemma 7.6 and the fact that $\Gamma$ is a graph, we have $(\gamma_{a^{-1}}\gamma_{a})\ast w[x]\eta=w[x]\eta$ and $(\gamma_{a}\gamma_{a^{-1}})\ast w[x]\eta=w[x]\eta$ for all $\eta\in\Lambda^{\omega}$. Moreover, by Lemma 7.5 for all $\xi\in\Lambda^{\omega}\setminus\mathcal{F}(\Lambda)$ we have $(\gamma_{a^{-1}}\gamma_{a})\ast\xi=\xi$ and $(\gamma_{a}\gamma_{a^{-1}})\ast\xi=\xi$. Therefore, we may conclude that each state $\gamma_{a}$ of $\mathcal{T}_{\Gamma}$ gives rise to a rational bijection $\gamma_{a}\ast:\Lambda^{\omega}\to\Lambda^{\omega}$ with inverse $\gamma_{a^{-1}}\ast$. Thus, by [4, Proposition 1.3], each state $\gamma_{a}$ is a rational homeomorphism of $\Lambda^{\omega}$. Let us prove that the subgroup $H(\Gamma)=\langle\gamma_{a}:a\in A\rangle$ of the rational group $\mathcal{R}_{\Lambda}$ of all the rational homeomorphisms of $\Lambda^{\omega}$ is isomorphic to the transtion group $\mathcal{G}(\Gamma)$. Consider a word $u=a_{1}\ldots a_{p}\in A^{*}$. By Lemma 7.6, and using a simple induction on the length $p$ of a word $u$ we have that the following equality
@@
8. Perturbing a graph: a structural result
Another natural question about these groups is how much small local changes in the graph affect the associated transition groups. In this section we investigate what happens to the transition group when we locally perturb a context-free graph $(\Gamma,x_{0})$. The idea is to consider a disk $D_{n}(x_{0})$ and to associate with the graph $\Lambda_{n}=\Gamma\setminus D_{n}(x_{0})$ a group called the boundary group $\mathrsfs{B}_{n}$ at level $n$, and then to give a structural result connecting this group to the transition group $G=\mathcal{G}(\Gamma)$.
Definition 8.1 (Boundary group).
Consider the following subset of $A^{*}$
[…]
We have the following key lemma.
Lemma 8.2.
Proof.
Consider the infinite walk $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, and let us denote by $x_{i}^{(j)}$ the vertices of such walk with $x=x_{0}^{(0)}$ and $x_{i}^{(j)}$ for $0\leq i\leq|u|$ are all the vertices between $x_{0}^{(j)}$ and $x_{0}^{(j+1)}$ so that $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ are the corresponding subwalks of $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots$. If $x\cdot u=x$ we are done, thus we may suppose $x\cdot u\neq x$. Note that every subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ intersects $S_{n}(x_{0})$. Indeed, if $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ were totally contained in $\Lambda_{n}$, then since $u\in\mathcal{O}_{n}$, we would have $x_{0}^{(j)}=x_{0}^{(j+1)}$, hence $x\cdot u=x$ when $j=0$, a contradiction. Now if we have that $x_{i}^{{(j)}}=x_{i}^{{(s)}}=v\in S_{n}(x_{0})$, for some $s\neq j$, then by the determinism of $\Gamma$ we have $x\cdot u^{k}=x$ for some $k$. Since the walk $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ is infinite and $S_{n}(x_{0})$ is a finite set. Either one of the two cases happens. To give the upper bound on $k$, we consider the worst case scenario: each subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ contains at most a vertex of $S_{n}(x_{0})$, therefore by the pigeonhole principle it is easy to see that condition $x_{i}^{{(j)}}=x_{i}^{{(s)}}$ occurs for some $j,s\leq\left|S_{n}(x_{0})\right|\,\left|u\right|$. ∎
Corollary 8.3.
For all $n\geq 1$, the boundary group $\mathrsfs{B}_{n}$ at level $n$ is a torsion normal subgroup of $G$ where each element $g$ has order $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$ (where $|\cdot|$ is the group metric with respect to $A$).
In view of the structural result contained in Proposition 6.2 an interesting case is when the graph $\Lambda_{n}$ can be obtained by a quasi-transitive graph by cutting out the disk $D_{n}(x_{0})$, more precisely we define the following class of graphs.
Definition 8.4.
A graph $(\Gamma,x_{0})$ is called locally-quasi-transitive if there is a quasi-transitive graph $(\Theta,y_{0})$ and an integer $n\geq 1$ such that $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$.
Lemma 8.5.
$\mathcal{O}_{n}=\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$
Proof.
The equality $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ follows by Proposition 2.5 as two vertices in the same orbits have the same associate language. Since $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ are words that label a circuit at every point, we clearly have the inclusion $\bigcap_{p\in R}L(\Theta,p)\subseteq\mathcal{O}_{n}$. Conversely, take a word $u\in\mathcal{O}_{n}$, for each $p_{i}\in R$ we claim that $u$ labels a circuit at $p_{i}$. Indeed, since the orbit $O_{i}$ is infinite, we may find a vertex $p_{i}^{\prime}\in O_{i}$ that is at sufficiently large distance from the disk $D_{n}(x_{0})$ such that $p_{i}^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}z$ is totally contained in $\Lambda_{n}$, hence by definition $z=p_{i}^{\prime}$ and therefore $u$ labels also a circuit at $p_{i}$. Therefore, $u\in\bigcap_{p\in R}L(\Theta,p)$ showing the other inclusion. ∎
In case the graph $(\Gamma,x_{0})$ is locally-quasi-transitive where the associated quasi-transitive graph $\Theta$ has all the orbits that are infinite, we say that $(\Gamma,x_{0})$ is locally-quasi-transitive with infinite orbits. We are now in position to prove the following structural theorem.
Theorem 8.6.
Let $(\Gamma,x_{0})$ be a context-free graph that is also locally-quasi-transitive with infinite orbits and let $G=\mathcal{G}(\Gamma)$ be its transition group. Then, there is a short exact sequence
Proof.
Let $\Theta$ be the graph which is quasi-transitive with infinite orbits, such that there exists an integer $n\geq 1$ with $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$. Since there are finitely many end-cone types with frontier vertices contained inside the disk of radius $n$, we have that $\Gamma$ and $\Theta$ differ from finitely many end-cones up to end-isomorphism. Hence, $\Theta$ is also context-free. Therefore, by Proposition 6.2 we have that $H=\mathcal{G}(\Theta)$ is virtually a finitely generated subgroup of the direct product of free groups. Now, by Lemma 8.5 we have $\mathcal{O}_{n}=\mathcal{L}(\Theta)$, hence since $\mathcal{L}(\Gamma)\subseteq\mathcal{O}_{n}=\mathcal{L}(\Theta)$, there is a well-defined epimorphism $\psi:\mathcal{G}(\Gamma)\rightarrow\mathcal{G}(\Theta)$, with $\ker(\psi)=\pi(\mathcal{O}_{n})=\mathrsfs{B}_{n}$ ($\pi$ is the natural map from $A^{*}$ to the group $\mathcal{G}(\Gamma)$). From which we have the short exact sequence in the statement and the fact that $H\simeq G/_{\mathrsfs{B}_{n}}$. Moreover, by Corollary 8.3, we have that the boundary group $\mathrsfs{B}_{n}$ is a torsion subgroup of $\mathcal{G}(\Gamma)$ where the order of each element $g$ is bounded above by $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$. ∎
It is worth mentioning that in the settings of the previous proposition, the boundary group $\mathrsfs{B}_{n}$ is generated by a poly-context-free language. This setting is going to be a key ingredient in Section 10 were we will address the second point of Question 6.5.
9. Free product of context-free graphs